OpenStudy (anonymous):

–3(5 – 15n) = 15(3n – 2) A. 30 B. –15 C. no solution D. all real numbers

OpenStudy (ryan123345):



OpenStudy (anonymous):

45 – 4r = 45 – 4r A. 1 B. 4 C. no solution D. all real numbers

OpenStudy (ryan123345):

D

OpenStudy (anonymous):

Name the property for the given statement. (4 – 6)(3) = (4)(3) – (6)(3) A. addition property of equality B. distributive property C. multiplication property of equality D. subtraction property of equality